Blossom Company bought a machine on January 1, 2017. The machine cost $140000 and had an expected salvage value of $24000. The life of the machine was estimated to be 5 years. The company uses the straight-line method of depreciation. The book value of the machine at the beginning of the third year would be
Answer #1

Answer

> Book Value at the beginning of third year = Original Cost - Accumulated Depreciation of 2 years.

>Working

A Original cost $140,000
B Salvage value $24,000
C = A - B Depreciable base $116,000
D Life (in years)                                   5
E = C/D Straight Line Depreciation annual $23,200
Depreciation expense:
Year 1 $23,200
Year 2 $23,200
F Total Accumulated Depreciation for 2 years $46,400
G = A - F Book Value at the beginning of the third year $93,600 = Answer

>Correct Answer = $ 93,600